A steamboat went 8 miles upstream in 1 hour. The return trip took only 30 minutes. Assume that the speed of the current and the direction were constant during both parts of the trip. Find the speed of the boat in still water and the speed of the current.
Please help me figure out what 2 equations I would use and then I believe I can get the answer.
Thanks in advance for your help!

Answer by stanbon(75887)

A steamboat went 8 miles upstream in 1 hour.
The return trip took only 30 minutes.
Assume that the speed of the current and the direction were constant during both parts of the trip.
Find the speed of the boat in still water and the speed of the current.
-----------------------------
Upstream DATA:
distance = 8 miles ; time = 1 hr. ; rate = d/t = 8 mph
-------------------------------
Downstream DATA:
distance = 8 miles ; time = 1/2 hr. ; rate = d/t = 16 mph
-------------------------------
Equations:
boat - current = 8 mph
boat + current = 16 mph
-----------------------------
2*boat = 24 mph
boat = 12 mph (rate in still water)
--------------------------
Substitute into b + c = 16 to solve for current speed
12 + c = 16
c = 4 mph (current speed)
============================
Cheers,
Stan H.